# Multi-Amplifier Sensing Charge-coupled Devices for Next Generation Spectroscopy

### v1

@@ 5.1.2 Inter-amplifier charge transfer efficiency
To measure the low-signal ( electrons) inter-amplifier CTE (ACTE), we expose the CCD to X-rays. After overscan correction, the location of each X-ray event is identified by a standard source extraction algorithm, where aperture photometry is performed to obtain the conversion gain. The gain is computed for each MAS amplifier and each image output was converted to electrons from ADU to compensate for any gain variations between amplifiers. The peak pixel value in electrons and its corresponding coordinate for a cluster of pixels affected by an X-ray photon in the first MAS amplifier is compared against the number of electrons at this same pixel location in subsequent serial amplifiers. If charge transfer efficiency is perfect, the pixels should have an identical number of electrons at each MAS output stage. As Fig. 10 demonstrates, deviations from a unity slope is indicative of imperfect charge transfer between amplifiers. After correcting for the respective gains of the amplifiers, we find that the inter-amplifier efficiency between the first and last amplifiers is below the 2000 electron signal level, the energy range where most of the X-ray induced charges are concentrated. This was computed from measuring several X-ray exposures and averaging the result; the relatively high scatter in the measured ACTE value when repeating this method for many images drives the elevated uncertainty, motivating us to pursue a complementary approach to more precisely quantify the ACTE effect.
@@
This secondary method we have developed to quantify the ACTE involves linearly modeling charge transfer efficiency between the peak X-ray flux pixel in the first output stage with the adjacent pixel in the direction opposite to the serial shifts. This simple model, parameterized by the charge transfer inefficiency (CTI), is described by
@@
where is the signal, in number of electrons, at pixel coordinate of amplifier , and denotes the pixel with the highest flux from a given X-ray hit. In this model, the number of electrons in the pixel next to the X-ray peak pixel opposite to the serial shift direction (i.e., trailing pixel) is expressed as . This quantity is the sum of two components to first order: the number of electrons in this pixel as measured in the first amplifier (i.e., amplifier 0) with some loss due to non-zero CTI and the number of electrons lost from the peak X-ray pixel in the first amplifier due to non-zero CTI. A CTI value can be calculated for each X-ray hit on the image, and the distribution of these values can be fitted with a Gaussian to statistically estimate the ACTE as exhibited in Fig. 11. Using our lowest noise voltage configuration, we find that our lower limit for the inter-amplifier charge transfer efficiency between the first and last MAS amplifiers is by this statistical approach. This is in agreement with the simpler first approach illustrated by Fig. 10 to within uncertainties.
@@
However, X-rays are limited in energy range to the absorption lines of the 55Fe source, so probing for transfer problems at higher signal levels at the level and above requires an alternate approach. To reach higher illumination levels, the LED projector was triggered as a virtual shutter for a specified exposure time during which the CCD is in expose mode before the CCD is read out. The LED was connected externally to an external controller that connected to the synchronization lines of the Hydra and included options for adjusting the LED intensity.
@@
Our first approach is based loosely on the extended pixel edge response (EPER) method, which is a standard horizontal CTI measurement technique for conventional CCDs relying on flat field images to determine the amount of deferred charge in the overscan region. However, in addition to charge transfer in the inter-amplifier pixels, since charge enters and is removed from the charge sensing areas multiple times in the MAS architecture, the traditional EPER formulation is inadequate for analyzing the potential charge transfer loss and deferral between one output stage and the next. Thus, we follow Eq. 4 to compute the inter-amplifier CTI given the signal measured in the first overscan column between the first and last amplifiers. Applying this technique with EPER-like data exploits the sharp edge between the active and overscan regions. Because the overscan region is not expected to contain incident signal if the CTI is ideally zero, the sharp edge provides a sensitive probe of charge in the overscan area after shifts through multiple sense nodes that is independent of the surface uniformity of the illumination from imperfect optics.
@@
Illumination at varying signal levels was obtained by increasing the exposure time in increments of 250 ms up to 28 s in a linear ramp, with exposures taken at each exposure time. Using a reduced ADC gain, this ramp extends up to e- in incident signal level. The gain was again calibrated at this ADC configuration for each amplifier and all images were bias-subtracted with a master bias for each amplifier. The horizontal profiles are shown in Fig. 13, where the signal average of each column of pixels up to the vertical overscan are compared at various exposure times. The last active column is 538 and like typical LBNL -channel CCDs contains more charge than other pixels in the flat field. This effect may arise because charge from the flat field diffuses from material outside of the array at the edges (Janesick et al., 1989) and because fringing fields result in pixel size variations at the edges (Plazas et al., 2014). If we consider the first amplifier, as the signal level increases with the exposure time, at a certain point of high signal ( e-) the charge begins overflowing into the overscan region, leading to catastrophic charge loss in the active pixels which rapidly degrades as the active area completely saturates close to e- where the horizontal profiles develop the characteristic flat top. This is an expected effect present in each amplifier that reflects the physical pixel full well given our operating parameters. However, the horizontal profile for amplifier 16 shows that the charge overflow into the overscan pixels occurs earlier compared to the first amplifier, pointing to an inter-amplifier CTI effect that dominates the overall charge losses along the extended MAS register even before the pixel full well capacity is reached.
@@
To diagnose the amount of charge where the inter-amplifier charge transfer efficiency breaks down, we apply Eq. 4 for the mean of the final active column as the “peak pixel” and the first overscan column as its “trailing pixel.” This is evaluated at the first amplifier stage in relation to each of the subsequent stages from 5,000 to 80,000 e-. The results for this ACTE analysis is summarized in Fig. 14, where the measured ACTE for each output stage relative to amplifier 0 is plotted, showing the signal regime at which the amplifier-to-amplifier transfer drops. The inter-amplifier CTI between amplifier 0 and 1 is consistent with unity until above 53,000 e- where ACTE degrades swiftly. However for amplifier 15, the ACTE has a slightly steeper slope as seen in the inset of Fig. 14 and rapidly drops starting around 50,000 e-. This initial decay leads to a rapid turnover starting just under 55,000 e- where charge transfer between amplifiers becomes unreliable with increasing blooming of charge into the overscan columns. For scientific requirements where conventional CTI , this imposes a strict upper limit on the effective dynamic range of this MAS device with our operating parameters at e- for reading out through all 16 channels. This demonstrates that the amplifier-to-amplifier transfer efficiency breaks down well before the actual physical well depth of the pixels given our clock voltages, and thus acts as an “effective full well.” This behavior where the ACTE rather than the physical pixel well depth imposes an effective full well is analogous to how the output-amplifier voltage swing limits the full well of conventional CCDs (Bebek et al., 2017).
@@
Another method explored for this analysis involved projecting a pattern of a grid onto the detector surface. At the front of the LED, a mask lens with a grid pattern was inserted and manually aligned to the CCD rows and columns. The grid was focused by minimizing the width of the illuminated column that the LED projects onto the active area. The grid pattern mask ensures that the illumination level is consistent in a given column or row of the grid. As with flat fields, different signal levels in the illuminated areas of the grid could be obtained by varying the exposure time. To measure the CTE through the full dynamic range up to the full well capacity, we took images in a ramp of increasing exposure times up to 30 s at a low gain ADC setting.
@@ 5.2 Full well capacity
Our measurements with the inter-amplifier charge transfer efficiency indicate that there is an “effective full well” in the MAS architecture bounded by non-linearities at high signal that is distinct from the physical well depth of the pixel given the clock voltages. These non-linearities result from the highly inefficient charge transfers between amplifier sense nodes in the MAS register that dominate all other effects starting at the 50,000 e- signal level. This is specifically the limit contributed by amplifier 15, the final output stage, and thus can be interpreted as an effective charge level constraint for the device. Because this effective full well is not set by the pixel well depth but by the PS-H1 clock voltages that dictate ACTE, it is another parameter that should be optimized on, akin to how the sense-node conversion gain impacts the full well of conventional CCDs.
@@ 5.4 Crosstalk
Multiple channels measure charge simultaneously in the MAS CCD, introducing a risk of electronic crosstalk between channels that generates ghost images that can contaminate neighboring channels. One indication of crosstalk we observed especially in flat fields was the slightly different baseline levels within the overscan and prescan regions of the image. In column-averaged horizontal profiles of the prescan and overscan regions, after the sharp drop-off in signal from the active region, the first 16 pixels into the overscan form a “step” before dropping to a stable baseline. This 16-pixel wide step corresponds to the number of inter-amplifier pixels separating the output stages, and is characteristic of crosstalk generated by amplifiers that are reading pixels at the same time for pixels physically ahead and behind horizontally. The signal difference between this 16-pixel step and the overscan baseline scales linearly with the signal in the final active column until inter-amplifier CTI losses cause blooming into the overscan itself. Using this feature to estimate the level of this ghosting signature present in each channel, we found the upper limit of this crosstalk to be of the signal in the final active column in the amplifiers 4-8, in the middle of the chain. The amplifiers with lower levels of ghosting were 0-3 and 9-15, where the crosstalk amplitude was measured to be , comparable to common requirements for multi-output conventional CCDs (e.g., O’Connor, 2015).
@@ 6 Upcoming plans
For photon-limited observations required for the more aggressive space imaging application, single-photon resolution is essential and necessitates a combination of further reductions in noise per amplifier, additional MAS output stages, and taking advantage of the ability to “skip” with the floating-gate amplifier. In the photon-counting regime, systematic noise backgrounds normally hidden underneath read noise in silicon detectors become dominant, imposing a potential detection limit that must be characterized. When the magnitudes of the carriers created from thermal leakage (dark) current and clock-induced charge are comparable to incident optical signals, the photons are to be detected become indistinguishable from the background generated from the detector itself. While amplifier thermal leakage is typically managed by cooling the detector, clock-induced charge spuriously generated during readout is a well-documented but not yet fully-understood single-electron event mechanism that single-photon resolving Skipper and electron-multiplying CCDs are sensitive to detecting (e.g., Janesick, 2001; Barak et al., 2022; Wilkins et al., 2014). Because the MAS CCD has additional clocks owing from inter-amplifier pixels in its extended output register, investigating the nature of spurious charge in the photon-counting regime is especially important, and our next phase of testing enables output gate clocking for “skipping” capability to extensively quantify its dependence on clock amplitude, edge shape, and pixel rate. Re-sampling charge to reach photon-counting resolution also provides the additional benefit of direct gain self-calibration and very low-signal linearity measurements during potential daytime detector calibrations in survey operations. New readout schemes that benefit from multiple amplifiers will also be studied. To accommodate these functionalities with a large format MAS CCD, a 64-channel readout electronics system has been designed where these tests will be carried out.
@@ 7 Conclusions
The progress presented here is the beginning of an extensive effort to fully optimize a MAS CCD as a candidate detector for future spectroscopic experiments and space-based imaging such as HWO. Similar results with respect to noise and inter-amplifier charge transfer efficiency have also been found with a 16 channel MAS CCD operating with LTA readout electronics (Botti et al., 2024; Lapi et al., 2024). With a fundamental proof-of-concept presented here, upcoming large format, high channel count designs have been pushed forward for prototyping, and new readout electronics solutions that accommodate high channel counts including ASICs are being explored.

### v2

@@ 5.1.2 Inter-amplifier charge transfer efficiency
To measure the low-signal ( electrons) inter-amplifier CTE (ACTE), we expose the CCD to X-rays. After overscan correction, the location of each X-ray event is identified by a standard source extraction algorithm (Barbary, 2016), where aperture photometry is performed to obtain the conversion gain. This process for this was described in 4.1. The gain is computed for each MAS amplifier and each image output was converted to electrons from ADU to compensate for any gain variations between amplifiers. The peak pixel value in electrons and its corresponding coordinate for a cluster of pixels affected by an X-ray photon in the first MAS amplifier is compared against the number of electrons at this same pixel location in subsequent serial amplifiers. If charge transfer efficiency is perfect, the pixels should have an identical number of electrons at each MAS output stage. As Fig. 10 demonstrates, deviations from a unity slope is indicative of imperfect charge transfer between amplifiers. After correcting for the respective gains of the amplifiers, we find that the inter-amplifier efficiency between the first and last amplifiers is below the 2000 electron signal level, the energy range where most of the X-ray induced charges are concentrated. This was computed from measuring several X-ray exposures and averaging the result; the relatively high scatter in the measured ACTE value when repeating this method for many images drives the elevated uncertainty, motivating us to pursue a complementary approach to more precisely quantify the ACTE effect.
@@
This secondary method we have developed to quantify the ACTE involves linearly modeling charge transfer efficiency between the peak X-ray flux pixel in the first output stage with the adjacent pixel in the direction opposite to the serial shifts. This simple model, parameterized by the charge transfer inefficiency, is described by
@@
where is the signal, in number of electrons, at pixel coordinate of amplifier , and denotes the pixel with the highest flux from a given X-ray hit. In this model, the number of electrons in the pixel next to the X-ray peak pixel opposite to the serial shift direction (i.e., trailing pixel) is expressed as . This quantity is the sum of two components to first order: the number of electrons in this pixel as measured in the first amplifier (i.e., amplifier 0) with some loss due to non-zero CTI and the number of electrons lost from the peak X-ray pixel in the first amplifier due to non-zero CTI. A CTI value can be calculated for each X-ray hit on the image, and the distribution of these values can be fitted with a Gaussian to statistically estimate the ACTE as exhibited in Fig. 11. Using our lowest noise voltage configuration, we find that our lower limit for the inter-amplifier charge transfer efficiency between the first and last MAS amplifiers is by this statistical approach. This is in agreement with the simpler first approach illustrated by Fig. 10 to within uncertainties. Since this is measured at the level of 0.3%, the effect of adding additional higher order terms is negligible and would not change our results.
@@
However, X-rays are limited in energy range to the absorption lines of the 55Fe source, so probing for transfer problems at higher signal levels at the level and above requires an alternate approach. To reach higher illumination levels, the LED projector was triggered by specifying an exposure time during which the LED would be on and the CCD would be in expose mode. After the exposure time elapsed, the LED would turn off and the CCD readout would begin. By synchronizing the exposure time with the LED uptime, the LED acts as a virtual shutter with no mechanical shutter needed in our setup. The LED was connected externally to an external controller that connected to the synchronization lines of the Hydra and included options for adjusting the LED intensity.
@@
Our first approach is based loosely on the extended pixel edge response (EPER) method, which is a standard horizontal CTI measurement technique for conventional CCDs relying on flat field images to determine the amount of deferred charge in the overscan region (Janesick et al., 1988). However, in addition to charge transfer in the inter-amplifier pixels, since charge enters and is removed from the charge sensing areas multiple times in the MAS architecture, the traditional EPER formulation is inadequate for analyzing the potential charge transfer loss and deferral between one output stage and the next. Thus, we follow Eq. 4 to compute the inter-amplifier CTI given the signal measured in the first overscan column between the first and last amplifiers. Applying this technique with EPER-like data exploits the sharp edge between the active and overscan regions. Because the overscan region is not expected to contain incident signal if the CTI is ideally zero, the sharp edge provides a sensitive probe of charge in the overscan area after shifts through multiple sense nodes that is independent of the surface uniformity of the illumination from imperfect optics.
@@
Illumination at varying signal levels was obtained by increasing the exposure time in increments of 250 ms up to 28 s in a linear ramp, with exposures taken at each exposure time. Using a reduced ADC gain, this ramp extends up to e- in incident signal level. The gain was again calibrated at this ADC configuration for each amplifier and all images were bias-subtracted with a master bias for each amplifier. The master bias was obtained by averaging a stack of 40 bias frames on a per-pixel basis. The horizontal profiles are shown in Fig. 13, where the signal average of each column of pixels up to the vertical overscan are compared at various exposure times. The last active column is 538 and like typical LBNL -channel CCDs contains more charge than other pixels in the flat field. This effect may arise because charge from the flat field diffuses from material outside of the array at the edges (Janesick et al., 1989) and because fringing fields result in pixel size variations at the edges (Plazas et al., 2014). If we consider the first amplifier, as the signal level increases with the exposure time and saturation approaches ( e-) the charge begins overflowing into the overscan region. As the pixels in the active area completely saturate, the horizontal profile over the width of the CCD develops a characteristic flat top. This is an expected effect present in each amplifier that reflects the physical pixel full well given our operating parameters. However, the horizontal profile for amplifier 15 shows that the charge overflow into the overscan pixels occurs earlier compared to the first amplifier, pointing to an inter-amplifier CTI effect that dominates the overall charge losses along the extended MAS register even before the pixel full well capacity is reached.
@@
To diagnose the amount of charge where the inter-amplifier charge transfer efficiency breaks down, we apply Eq. 4 for the mean of the final active column as the “peak pixel” and the first overscan column as its “trailing pixel.” This is evaluated at the first amplifier stage in relation to each of the subsequent stages from 5,000 to 80,000 e-. The results for this ACTE analysis is summarized in Fig. 14, where the measured ACTE for each output stage relative to amplifier 0 is plotted, showing the signal regime at which the amplifier-to-amplifier transfer drops. The inter-amplifier CTI between amplifier 0 and 1 is consistent with unity until above 53,000 e- where ACTE degrades swiftly. However for amplifier 15, the ACTE has a slightly steeper slope as seen in the inset of Fig. 14 and rapidly drops starting around 50,000 e-. This initial decay leads to a rapid turnover starting just under 55,000 e- where charge transfer between amplifiers becomes unreliable with increasing blooming of charge into the overscan columns. For scientific requirements where conventional CTI , this imposes a strict upper limit on the effective dynamic range of this MAS device with our operating parameters at e- for reading out through all 16 channels. This demonstrates that the amplifier-to-amplifier transfer efficiency breaks down well before the actual physical well depth of the pixels given our clock voltages, and thus acts as an “effective full well.” This behavior where the ACTE rather than the physical pixel well depth imposes an effective full well is analogous to how the output-amplifier voltage swing limits the full well of conventional CCDs (Bebek et al., 2017). We note that although the inter-amplifier CTI at amplifier 10 is larger than in Fig. 12, the requirements for pixel-to-pixel CTE are different from the inter-amplifier CTE. This is because the nominal CTI requirement is an average over thousands of pixel shifts, while the ACTE is measured over 16 amplifiers and the maximum combined effect is a 0.3% inefficiency. For a CTI standard over 1000 shifts, this translates to an integrated CTI requirement of , and this is met even with the amplifier 10 defect.
@@
Another method explored for this analysis involved projecting a pattern of a grid onto the detector surface. At the front of the LED, a mask lens with a grid pattern was inserted and manually aligned to the CCD rows and columns. Alignment was done by rotating the LED projector with the mask lens attached, taking a sample image, and evaluating whether the center lines of the illuminated grid columns and/or rows had flat profiles, since in the illuminated grid regions the signal level should be flat. The grid was focused by minimizing the width of the illuminated column that the LED projects onto the active area. The grid pattern mask ensures that the illumination level is consistent in a given column or row of the grid. As with flat fields, different signal levels in the illuminated areas of the grid could be obtained by varying the exposure time. To measure the CTE through the full dynamic range up to the full well capacity, we took images in a ramp of increasing exposure times up to 30 s at a low gain ADC setting.
@@ 5.2 Full well capacity
Our measurements with the inter-amplifier charge transfer efficiency indicate that there is an “effective full well” in the MAS architecture bounded by non-linearities at high signal that is distinct from the physical well depth of the pixel given the clock voltages. These non-linearities result from the highly inefficient charge transfers between amplifier sense nodes in the MAS register that dominate all other effects starting at the 50,000 e- signal level. This is specifically the limit contributed by amplifier 15, the final output stage, and thus can be interpreted as an effective charge level constraint for the device. Because this effective full well is not set by the pixel well depth but by the PS-H1 clock voltages that dictate ACTE, it is another parameter that should be optimized on, akin to how the sense-node conversion gain impacts the full well of conventional CCDs (e.g., Fossum & Hondongwa, 2014; Bebek et al., 2017).
@@ 5.4 Crosstalk
Multiple channels measure charge simultaneously in the MAS CCD, introducing an elevated risk of electronic crosstalk between channels that generates ghost images that can contaminate neighboring channels. One indication of crosstalk we observed especially in flat fields was the slightly different baseline levels within the overscan and prescan regions of the image. In column-averaged horizontal profiles of the prescan and overscan regions, after the sharp drop-off in signal from the active region, the first 16 pixels into the overscan form a “step” before dropping to a stable baseline. This 16-pixel wide step corresponds to the number of inter-amplifier pixels separating the output stages, and is characteristic of crosstalk generated by amplifiers that are reading pixels at the same time for pixels physically ahead and behind horizontally. The signal difference between this 16-pixel step and the overscan baseline scales linearly with the signal in the final active column until inter-amplifier CTI losses cause blooming into the overscan itself. Using this feature to estimate the level of this ghosting signature present in each channel, we found the upper limit of this crosstalk to be of the signal in the final active column in the amplifiers 4-8, in the middle of the chain. The amplifiers with lower levels of ghosting were 0-3 and 9-14, where the crosstalk amplitude was measured to be , comparable to common requirements for multi-output conventional CCDs (e.g., Freyhammer et al., 2001; O’Connor, 2015).
@@ 6 Upcoming plans
For photon-limited observations required for the more aggressive space imaging application, single-photon resolution is essential and necessitates a combination of further reductions in noise per amplifier, additional MAS output stages, and taking advantage of the ability to “skip” with the floating-gate amplifier. In the photon-counting regime, systematic noise backgrounds normally hidden underneath read noise in silicon detectors become dominant, imposing a potential detection limit that must be characterized. When the magnitudes of the carriers created from thermal leakage (dark) current and clock-induced charge are comparable to incident optical signals, the photons are to be detected become indistinguishable from the background generated from the detector itself. While amplifier thermal leakage is typically managed by cooling the detector, clock-induced charge spuriously generated during readout is a well-documented but not yet fully-understood single-electron event mechanism that single-photon resolving Skipper and electron-multiplying CCDs are sensitive to detecting (e.g., Janesick, 2001; Barak et al., 2022; Wilkins et al., 2014). Because the MAS CCD has additional clocks owing from inter-amplifier pixels in its extended output register, investigating the nature of spurious charge in the photon-counting regime is especially important, and our next phase of testing enables output gate clocking for “skipping” capability to extensively quantify its dependence on clock amplitude, edge shape, and pixel rate. Re-sampling charge to reach photon-counting resolution also provides the additional benefit of direct gain self-calibration and very low-signal linearity measurements during potential daytime detector calibrations in survey operations. New readout schemes that benefit from multiple amplifiers will also be studied. To accommodate these functionalities with a large format MAS CCD, a 64-channel readout electronics system is being designed that will be used to carry out these tests.
@@ 7 Conclusions
The progress presented here is the beginning of an extensive effort to fully optimize a MAS CCD as a candidate detector for future spectroscopic experiments and space-based imaging such as HWO. Similar results with respect to noise and inter-amplifier charge transfer efficiency have also been found with a 16 channel MAS CCD operating with LTA readout electronics (Botti et al., 2024; Lapi et al., 2024). With a fundamental proof-of-concept presented here, upcoming large format, high channel count designs have been pushed forward for prototyping, and new readout electronics solutions that accommodate high channel counts including ASICs are being explored (e.g., Loose et al., 2018; Bessia et al., 2023).